How do you put 4y-2x=4 into y=mxtb?
Answer by crabbyoldgamer(13)

You use algebra.
4y-2x=4 : add 2x to both sides:
4y = 2x + 4: divide both sides by 4.
y = (1/2)x + 1: m = 1/2, b = 1